# Understanding Emotion in Discourse: Recognition Insights and Linguistic Patterns for Generation

### v1

@@ 3.4 Training and Implementation Details
Our classifier consists of a two-layer MLP (for K=0) or a single-layer unidirectional LSTM (for K>0). We use fixed hyperparameters across all experiments: learning rate of 1e-3, hidden dimension of 256, dropout rate of 0.3, and batch size of 64. Training employs Adam optimizer with early stopping based on validation loss (patience of 60 epochs for utterance-level, 20 epochs for turn-level experiments).
All experiments use 10 random seeds {42, 43, …, 51} with results reported as mean $\pm$ standard deviation. Implementation uses PyTorch 1.13 with experiments conducted on NVIDIA A100 GPUs via Saturn Cloud.
4 Experiments and Analysis
@@ 4.1 Encoder Selection
We evaluated three pre-trained encoders (Table 1). Sentence-RoBERTa achieves the best performance (65.29%), likely due to its NLI fine-tuning which implicitly captures affective reasoning. All subsequent experiments use Sentence-RoBERTa.
@@ 4.2 Main Results
Table 2 presents our main results comparing flat and hierarchical encoding strategies. A striking pattern emerges: at utterance-level (K=0), hierarchical encoding outperforms flat encoding (+3.17%p for 4-way, $p<.01$; +1.80%p for 6-way, $p<.05$), suggesting that intra-utterance sentence structure aids emotion recognition. However, when conversational context is incorporated, flat encoding achieves the best performance (82.69% for 4-way, 67.07% for 6-way), though the difference from hierarchical encoding is not statistically significant ($p=.427$ for 4-way, $p=.078$ for 6-way). This reversal suggests that turn-level context subsumes the structural information captured by hierarchical encoding. We note that variations in layer selection (last vs. avg_last4) and pooling methods showed no significant differences (14 comparisons, all $p>.08$; see Appendix), justifying our reporting of only the best configurations.
Table 2: Main results on IEMOCAP (10 seeds). $K$ denotes the number of preceding turns (past-only context). Each row shows the best configuration for that encoding strategy. Hierarchical encoding helps at utterance-level (K=0), but flat encoding wins with conversational context.
Task K Encoding Layer Pool Mean $\pm$ Std Min Max 95% CI
4-way 0 FLAT last wmean_pos 65.29 $\pm$ 1.17 64.10 67.31 [64.45, 66.12]
4-way 0 HIER avg_last4 wmean_pos_rev 68.46 $\pm$ 1.09 66.30 70.12 [67.68, 69.24]
4-way 132 FLAT avg_last4 mean 82.69 $\pm$ 0.50 81.73 83.43 [82.33, 83.05]
4-way 132 HIER avg_last4 mean 81.89 $\pm$ 0.41 81.22 82.60 [81.60, 82.18]
6-way 0 FLAT last wmean_pos 52.69 $\pm$ 1.04 50.61 54.51 [51.95, 53.43]
6-way 0 HIER avg_last4 wmean_pos_rev 54.49 $\pm$ 0.84 53.17 55.77 [53.88, 55.09]
6-way 101 FLAT avg_last4 wmean_pos 67.07 $\pm$ 0.69 65.75 68.04 [66.58, 67.57]
6-way 101 HIER avg_last4 mean 66.73 $\pm$ 0.87 65.53 68.68 [66.11, 67.36]
@@ 4.3 Emotion-Specific Context Effects
While prior work on variable-length context focuses on how to adaptively select context windows through speaker-aware modules (zhang2023emotion), we take a complementary approach: we systematically investigate what patterns emerge when varying context length and whether different emotions exhibit distinct context requirements.
To analyze emotion-specific effects, we compute per-class F1 scores at each context length from K=0 (utterance only) to K=200 (preceding turns). For each emotion, we report the context length at which F1 attains its maximum within the sweep range, and quantify improvement relative to $K=0$. We also determine the saturation point, defined as the minimum K at which 90% of maximum improvement is achieved. All analyses are conducted across 10 random seeds.
Our first finding is that performance saturates rapidly. As shown in Figure 3, 90% of maximum improvement is achieved with K=30 for 4-way classification and K=10 for 6-way classification. These saturation points correspond to a small fraction of typical dialogue history measured in turns. This suggests that the immediate preceding context carries most of the predictive information for emotion recognition, and sophisticated adaptive selection mechanisms may offer diminishing returns for typical utterances.
Our second finding reveals a dissociation between saturation timing and improvement magnitude. While all emotions reach saturation at similar context lengths (Kruskal-Wallis $H=0.52$, $p=0.91$; mean saturation K ranges from 43 to 61 across emotions), the magnitude of improvement differs dramatically across emotions (one-way ANOVA $F=136.80$, $p<0.0001$). Table 3 presents the full breakdown. For 4-way classification, Sad shows the largest improvement (+22.31 percentage points), followed by Happy (+15.82%p), Neutral (+10.77%p), and Angry (+8.34%p). This ordering is remarkably consistent in 6-way classification, where Sad again benefits most (+21.58%p) and Angry benefits least (+9.43%p), with Frustrated showing similarly low improvement (+8.73%p).
The context length at which each emotion attains its peak performance varies. For 4-way classification, Angry peaks at $K=10$ (10 preceding turns), whereas Sad, Happy, and Neutral peak at K=130K=130140140 (130–140 preceding turns). In 6-way classification, Sad peaks at $K=200$, while Angry peaks at $K=60$. These patterns suggest that certain emotions are recognizable from immediate local cues, while others depend on accumulated conversational dynamics.
We discuss the theoretical implications of these findings, including why arousal alone cannot explain the observed patterns and what this reveals about the linguistic expression of different emotions, in Section 5.
Figure 3: Per-emotion F1 scores across context sizes (4-way classification). Angry saturates quickly at $K^{*}=10$ (10 preceding turns), while Sad, Happy, and Neutral require extended context ($K^{*}=130$140140 preceding turns). Shaded regions indicate $\pm 1$ standard deviation across 10 seeds.
[…]
4-way Classification
[…]
Sad 46.17 130 +22.31
Happy 58.53 140 +15.82
Neutral 43.92 140 +10.77
Angry 35.27 10 +8.34
6-way Classification
[…]
Sad 46.19 200 +21.58
Excited 51.68 90 +10.99
Neutral 43.23 100 +10.69
Happy 31.73 110 +9.68
Angry 41.11 60 +9.43
Frust. 46.66 90 +8.73
@@ 4.4 Ablation Studies
We conducted ablation studies on pooling strategy, external lexical knowledge, and layer selection.
For pooling strategy, we compared mean pooling, position-weighted pooling emphasizing utterance-final tokens (wmean_pos), and position-weighted pooling emphasizing utterance-initial tokens (wmean_pos_rev). No method achieved statistically significant superiority (Friedman test, $p>.08$). Position-weighted pooling yielded the best configurations at utterance-level (K=0), while mean pooling performed best with conversational context (K>0).
For external lexical knowledge, we integrated SenticNet, which provides 4-dimensional affective ratings (pleasantness, attention, sensitivity, aptitude). However, this yielded no improvement (4-way: 0.94-0.94%; 6-way: $\sim$0%), suggesting that pre-trained language model embeddings already capture sufficient affective semantics (see Appendix C for complete results across 36 configurations).
For layer selection, averaging the last four transformer layers (avg_last4) versus using only the final layer (last) showed no significant difference (paired $t$-tests, min $p=.244$).
@@ 4.5 Comparison with Prior Work
We compare our approach against prior text-only methods on IEMOCAP. A critical distinction among ERC methods is their temporal context access: bidirectional methods utilize both past and future utterances, while past-only (causal) methods access only preceding context. This distinction is crucial for real-time deployment scenarios where future utterances are unavailable.
@@
Table 4 shows that our past-only approach achieves the highest performance among all text-only methods, including those using bidirectional context. We outperform HFFN by +1.15%p and HCAM by +1.29%p, despite these methods having access to future context. This result demonstrates that our discourse-aware pooling and emotion-specific context modeling capture patterns that bidirectional architectures miss.
@@
Ours Past-only 82.69
HFFN (mai2019divide) Bidirectional 81.54
HCAM (dutta2024hcam) Bidirectional 81.4
CHFusion (majumder2018multimodal) Bidirectional 73.6
@@
Table 5 presents results on the more challenging 6-way task. Among past-only methods, DAG-ERC achieves 68.03%, while our mean performance across 10 seeds is 67.07%. While we report mean performance across 10 seeds for statistical rigor, our best run achieves 68.04%, marginally surpassing DAG-ERC’s reported 68.03%. More notably, our past-only approach outperforms six bidirectional methods, demonstrating that access to future context does not guarantee superior performance when discourse-aware features effectively capture emotional dynamics from past context alone.
The 2.42%p gap between EmoCaps and our method can be attributed to EmoCaps’ bidirectional context and emotion capsule architecture specifically designed for capturing emotional tendencies. Nevertheless, the competitive performance of our simpler, strictly causal approach suggests that carefully designed discourse features can partially compensate for the absence of future context.
Table 5: Comparison on IEMOCAP 6-way classification (text-only methods, excluding LLM-based approaches). Bold indicates best past-only result.
@@
EmoCaps (li2022emocaps) Bidirectional 69.49
DAG-ERC (shen2021directed) Past-only 68.03
Ours Past-only 67.07
SKAIG (li2021past) Bidirectional 66.96
DialogueCRN (hu2021dialoguecrn) Bidirectional 66.20
CoG-BART (li2022cogbart) Bidirectional 66.18
BiERU (li2022bieru) Bidirectional 64.59
HCAM (dutta2024hcam) Bidirectional 64.4
DialogueGCN (ghosal2019dialoguegcn) Bidirectional 64.18
@@ 5 Discussion
Our experiments yield state-of-the-art performance among past-only methods (82.69% for 4-way, 67.07% for 6-way) while revealing several interpretable patterns: hierarchical encoding helps only without context, emotions differ dramatically in context requirements (+8%p to +22%p), and external lexicons provide no benefit. Our discourse marker analysis of 5,286 occurrences uncovers emotion-specific positioning patterns. Below, we discuss four key implications of these findings.
5.1 Conversational Context Subsumes Hierarchical Structure
Our results reveal a striking interaction between encoding strategy and context availability. At utterance-level (K=0), hierarchical encoding outperforms flat encoding by +3.17%p (4-way) and +1.80%p (6-way), confirming that intra-utterance sentence structure provides useful signal for emotion recognition. However, this advantage disappears—and even reverses—when conversational context is incorporated: flat encoding achieves 82.69% versus 81.89% for hierarchical (4-way), and 67.07% versus 66.73% (6-way).
This pattern suggests that turn-level context subsumes the structural information captured by hierarchical encoding. When the model can access preceding utterances, it learns emotional dynamics from the conversation flow, rendering fine-grained sentence boundaries within a single utterance redundant. This finding has practical implications: simpler flat architectures suffice when sufficient conversational context is available, reducing computational overhead without sacrificing performance.
5.2 Do Utterance Peripheries Carry Emotional Information?
Although pooling strategies did not yield statistically significant differences (Friedman $p>.08$), the pattern of best configurations reveals a meaningful connection to discourse structure. At utterance-level (K=0), position-weighted pooling consistently outperformed mean pooling: wmean_pos (emphasizing final tokens) for flat encoding, wmean_pos_rev (emphasizing initial tokens) for hierarchical encoding. With conversational context (K>>0), this advantage disappeared and mean pooling sufficed.
This pattern aligns with our discourse marker analysis. We found a significant association between emotion and peripheral positioning ($\chi^{2}$ test, $p<.0001$, Cramér’s $V=0.062$). Critically, Sad utterances show reduced left-periphery usage (21.9%) compared to Neutral (31.7%), Happy (29.7%), and Angry (28.2%). Post-hoc comparisons confirm Sad differs significantly from all other emotions (all $p<.01$ with Bonferroni correction).
Left-periphery markers like “well” and “oh” signal floor-claiming and turn-management—functions associated with active discourse engagement. The reduced left-periphery usage in Sad utterances suggests that sadness manifests through diminished pragmatic signaling, consistent with the low-arousal, withdrawal-oriented nature of this emotion. When conversational context is unavailable, position-weighted pooling captures these peripheral cues; when context is available, inter-turn dynamics render this positional information redundant.
5.3 Why Does Sadness Require More Context?
Our emotion-specific analysis reveals that Sad benefits most from context (+22.31%p) while Angry benefits least (+8.34%p). An arousal-based explanation would predict that high-arousal emotions are expressed explicitly while low-arousal emotions require contextual inference. However, this account fails: Happy is high-arousal yet benefits substantially from context (+15.82%p), far more than Angry.
We propose that the key factor is not arousal but explicitness of linguistic markers. Angry utterances contain salient lexical cues—profanity, emphatic negation (“I can’t believe”), exclamatory expressions—that are recognizable without conversational history. Sad utterances, in contrast, often lack explicit markers. Expressions like “I see,” “Oh,” or “Yeah, I guess” are pragmatically ambiguous; their emotional valence emerges only from the preceding conversational trajectory.
This interpretation connects to our discourse marker findings: Sad shows reduced use of explicit pragmatic signals (left-periphery markers), forcing the model to rely on accumulated context for disambiguation. Happy presents an intermediate case—positive expressions can be ambiguous with sarcasm or polite neutrality, explaining why context helps despite high arousal. Angry, with its explicit linguistic markers, requires minimal contextual support.
5.4 Is the 6-way Taxonomy Linguistically Justified?
Our confusion matrix analysis raises fundamental questions about the appropriateness of IEMOCAP’s 6-way emotion taxonomy for text-based classification. Contrary to the intuition that emotions of similar valence should be more confusable, we find that Happy–Sad confusion (17.3% turn-level, 18.9% utterance-level) substantially exceeds Happy–Excited confusion (5.3% turn-level, 10.9% utterance-level). This suggests that text-based features do not align with the valence-based groupings assumed by the taxonomy.
The Happy–Excited confusion is also notably asymmetric: 20.6% of Excited samples are misclassified as Happy at utterance-level, while only 1.3% of Happy becomes Excited. This asymmetry suggests that Happy functions as a “catch-all” positive category, absorbing cases that lack the prosodic intensity markers distinguishing Excited. Without acoustic cues, the model defaults to the more frequent positive label.
These findings suggest that the 6-way taxonomy may conflate orthogonal dimensions—valence and arousal—into arbitrary categorical boundaries. For text-only ERC, researchers should consider dimensional models (valence-arousal space) as alternatives, hierarchical classification that separates valence before fine-grained categories, or reduced taxonomies that merge linguistically indistinguishable categories such as Happy and Excited.

### v2

@@ 3.4 Training and Implementation Details
Our classifier consists of a two-layer MLP (for K=0K{=}0) or a single-layer unidirectional LSTM (for K>0K{>}0). We use fixed hyperparameters across all experiments: learning rate $1\mathrm{e}{-3}$, hidden dimension 256, dropout 0.3, and batch size 64. Hyperparameters were selected once on the validation set and then held fixed across all ablations to ensure comparability.
Training employs Adam with early stopping on validation weighted F1 (patience 60 for utterance-level and 20 for turn-level experiments). [^4] We use validation weighted F1 to align the early-stopping criterion with the primary evaluation metric. All experiments use 10 random seeds $\{42,43,\ldots,51\}$ with results reported as mean $\pm$ standard deviation. Statistical comparisons across configurations use paired $t$-tests over seeds with Bonferroni correction where applicable. Implementation uses PyTorch 1.13, and experiments were conducted on NVIDIA A100 GPUs via Saturn Cloud.
@@ 3.5 Encoder Selection
We compared three pre-trained text encoders at utterance level (K=0K{=}0) on IEMOCAP 4-way classification (Table 1). Sentence-RoBERTa yields the highest mean weighted F1 (65.29%), which is plausibly attributable to its NLI fine-tuning that produces robust sentence-level semantics. We emphasize, however, that this ranking may be dataset-dependent: IEMOCAP consists of relatively well-structured, acted dyadic dialogues, and noisier, multi-party corpora (e.g., MELD) may favor different encoders. To isolate architectural effects under a fixed encoder, all subsequent experiments use Sentence-RoBERTa unless otherwise noted.
@@ 3.6 Main Results
We report two complementary views of performance. First, for a fair comparison across design choices, we evaluate each configuration at fixed context sizes $K$ (Table 2), where $K$ denotes the number of strictly preceding turns (past-only context). This controlled evaluation shows that conversational context dominates performance: most of the total gain is already realized with a short history (e.g., K=10K{=}10), and performance changes only marginally beyond K=30K{=}305050 depending on the task.
Within the same $K$, neither hierarchical encoding nor position-weighted pooling shows a consistent advantage. In particular, FLAT vs. HIER differences are not statistically significant in the fixed-$K$ setting (paired tests across seeds; p>0.9p{>}0.9), and pooling choices do not yield reliable improvements (all comparisons p>0.08p{>}0.08). These results suggest that once modest conversational history is provided, architectural sophistication at the utterance level contributes little relative to context.
Second, to characterize the performance–context trade-off and identify saturation behavior, we also conduct a full $K$-sweep analysis (Section 3.7). For completeness, we additionally report the best-performing configurations observed in the sweep (Table LABEL:tab:main-results); importantly, these are used for analysis rather than test-set model selection.
[…]
Task Type Pooling K=0K{=}0 K=10K{=}10 K=30K{=}30 K=50K{=}50
[…]
Task Type Pooling K=0K{=}0 K=10K{=}10 K=30K{=}30 K=50K{=}50
[…]
Task Type K=100K{=}100 K=200K{=}200
[…]
6-way HIER (wmean_pos_rev) 62.49 63.74
@@ 3.7 Emotion-Specific Context Effects
While prior work on variable-length context focuses on how to adaptively select context windows through speaker-aware modules (Zhang et al., 2023), we take a complementary approach: we systematically investigate what patterns emerge when varying context length and whether different emotions exhibit distinct context requirements.
To analyze emotion-specific effects, we compute per-class F1 scores at each context length from K=0K{=}0 (utterance only) to K=200K{=}200 (preceding turns). For each emotion, we report the context length at which F1 attains its maximum within the sweep range, and quantify improvement relative to K=0K{=}0. We also determine the saturation point, defined as the minimum $K$ at which 90% of the maximum improvement (within the sweep range) is achieved. All analyses are conducted across 10 random seeds.
Our first finding is that performance saturates rapidly in aggregate: a short history (e.g., K=10K{=}10) already recovers most of the eventual gain in weighted F1, and improvements beyond K=30K{=}305050 are modest for the average utterance (Table 2). This suggests that immediate preceding context carries the majority of predictive information for ERC under strictly causal access.
Our second finding reveals a dissociation between saturation timing and improvement magnitude at the emotion level. While emotions reach saturation at broadly similar context lengths (Kruskal–Wallis $H=0.52$, $p=0.91$), the magnitude of improvement differs substantially (one-way ANOVA $F=136.80$, $p<0.0001$). Sad benefits the most from context (+22%p), whereas Angry benefits the least (+8–9%p), and this ordering remains consistent across 4-way and 6-way settings.
Figure 3: Per-emotion F1 scores across context sizes (4-way classification). Angry peaks with relatively short context ($K^{*}\approx 10$), while Sad/Happy/Neutral peak at substantially larger $K$ values. Shaded regions indicate $\pm 1$ standard deviation across 10 seeds.
Finally, although aggregate performance plateaus quickly, the context length at which a given emotion attains its peak within the sweep range can be much larger. This long-tail behavior motivates reporting both fixed-$K$ comparisons (for fairness) and sweep-based analyses (for saturation characterization). We discuss theoretical implications in Section 4.
@@ 3.8 Ablation Studies
We conducted ablations over (i) pooling strategy, (ii) external lexical knowledge, and (iii) layer aggregation, using the same training protocol and 10-seed evaluation as in the main experiments.
Pooling strategy.
We compared mean pooling, position-weighted pooling emphasizing utterance-final tokens (wmean_pos), and position-weighted pooling emphasizing utterance-initial tokens (wmean_pos_rev). Across seeds, no pooling method was consistently superior (Friedman test, $p>.08$). In terms of best-performing configurations, position-weighted pooling more often appeared among the top runs at utterance-level (K=0K{=}0), whereas mean pooling was more frequently competitive when conversational context was included (K>0K{>}0); however, these differences were not statistically reliable.
External lexical knowledge.
We augmented utterance representations with SenticNet 7 (Cambria et al., 2024), which provides four affective dimensions (pleasantness, attention, sensitivity, aptitude). This augmentation did not improve performance and in some settings slightly reduced it (e.g., 4-way: $\approx-0.94$ F1 points; 6-way: near-zero change), suggesting that the pretrained encoder representations already capture the affective information that SenticNet provides. Full results across all configurations are reported in Appendix C.
Layer aggregation.
We compared averaging the last four transformer layers (avg_last4) against using only the final layer (last). We observed no significant differences (paired $t$-tests; minimum $p=.244$ across comparisons).
@@ 3.9 Comparison with Prior Work
We situate our results against prior text-only ERC methods on IEMOCAP. A key axis of comparison is temporal access: bidirectional models can use both past and future utterances, whereas past-only (causal) models are restricted to preceding context and are therefore deployable in real-time settings. Importantly, prior results are reported under heterogeneous protocols (e.g., context windows, architectures, and often single-run reporting), so the tables below provide a reference comparison rather than a strict head-to-head evaluation.
@@
Table 3 shows that our past-only model is competitive with, and in some cases exceeds, reported performance of representative text-only baselines, including several bidirectional systems. We report the mean across 10 seeds for statistical robustness, whereas most prior work reports a single number.
@@
Ours (mean over 10 seeds) Past-only 82.69
HFFN (Mai et al., 2019) Bidirectional 81.54
HCAM (Dutta and Ganapathy, 2024) Bidirectional 81.4
CHFusion (Majumder et al., 2018) Bidirectional 73.6
@@
On the 6-way task (Table 4), our past-only model achieves 67.07% weighted F1 (mean over 10 seeds), which is comparable to strong reported baselines and exceeds several bidirectional systems. We stress that differences should be interpreted cautiously due to protocol mismatch and the prevalence of single-run reporting in prior work.
Table 4: IEMOCAP 6-way (text-only, excluding LLM-based approaches). Our result is mean over 10 seeds; prior work numbers are reported. Bold indicates best reported past-only result among non-LLM baselines.
@@
EmoCaps (Li et al., 2022c) Bidirectional 69.49
DAG-ERC (Shen et al., 2021) Past-only 68.03
Ours (mean over 10 seeds) Past-only 67.07
SKAIG (Li et al., 2021) Bidirectional 66.96
DialogueCRN (Hu et al., 2021) Bidirectional 66.20
CoG-BART (Li et al., 2022a) Bidirectional 66.18
BiERU (Li et al., 2022b) Bidirectional 64.59
HCAM (Dutta and Ganapathy, 2024) Bidirectional 64.4
DialogueGCN (Ghosal et al., 2019) Bidirectional 64.18
@@ 4 Discussion
Our study delivers strong performance under a strictly causal (past-only) setting (82.69% for 4-way; 67.07% for 6-way) and yields several interpretable patterns. Across analyses, conversational context provides the dominant gain, emotion classes differ markedly in their reliance on context, and external lexicons (SenticNet) do not improve performance. In addition, our corpus-based analysis of 5,286 discourse-marker (DM) occurrences reveals small but reliable emotion-specific positional tendencies. Below, we discuss four implications.
4.1 Conversational Context Reduces the Marginal Utility of Hierarchical Structure
We observe an interaction between encoding strategy and context availability. In the best-performing configurations at K=0K{=}0, hierarchical encoding exceeds flat encoding (Table LABEL:tab:main-results), consistent with the intuition that intra-utterance sentence structure can provide additional cues when the model cannot consult surrounding turns.
However, once past conversational context is incorporated, this advantage diminishes: flat and hierarchical variants become closely matched, and the best-performing contextual models are not reliably separated by the choice of utterance encoder. This suggests that turn-level context can partly substitute for fine-grained intra-utterance structure, because emotional evidence is often distributed across adjacent turns (e.g., responses, clarifications, escalation/de-escalation). Practically, this supports using simpler flat encoders in causal, real-time deployments when sufficient past context is available.
4.2 Do Utterance Peripheries Carry Emotion-Relevant Signal?
Pooling variants did not yield statistically significant differences in our multi-seed comparisons (Friedman $p>.08$), yet we consistently observed that position-weighted pooling is competitive at K=0K{=}0, while simple mean pooling is sufficient once context is introduced. We interpret this as a weak but suggestive pattern rather than a definitive effect.
This interpretation is compatible with our discourse-marker analysis. We find a significant association between emotion and DM periphery category ($\chi^{2}$, $p<.0001$), although the effect size is small (Cramér’s $V=0.062$). Notably, Sad utterances show reduced left-periphery usage (21.9%) compared to Neutral (31.7%), Happy (29.7%), and Angry (28.2%), with post-hoc tests indicating that Sad differs from the other emotions (Bonferroni-corrected, all $p<.01$). One plausible account is that left-periphery markers (e.g., well, oh) often participate in turn-management and stance negotiation; their reduced use may reflect more muted pragmatic signaling in Sad speech. Importantly, we treat this as a linguistic hypothesis supported by corpus statistics, not as a proven mechanism of the pooling variants.
4.3 Why Does Sadness Benefit Most from Context?
Emotion-specific analysis shows that Sad gains the most from added context (+22.31%p), whereas Angry gains the least (+8.34%p). A pure arousal-based account is insufficient, because Happy (often high-arousal) also benefits substantially from context (+15.82%p). Instead, our results suggest that what matters is the availability of explicit lexical/pragmatic cues in the target utterance.
Angry turns often contain salient lexical signals (e.g., emphatic negation, direct accusations, profanity) that are informative even without prior history. By contrast, Sad turns can be lexically understated and pragmatically ambiguous (e.g., I see, yeah, I guess), making their emotional interpretation contingent on the preceding trajectory. This aligns with the DM finding that Sad shows reduced left-periphery marking, potentially reducing overt discourse-management signals and increasing reliance on conversational history for disambiguation. More broadly, this supports the view that causal ERC should be evaluated not only by overall F1 but also by how different emotions depend on context.
4.4 Is the 6-way Taxonomy Well-Identified in Text-Only ERC?
Our confusion patterns raise questions about how well the 6-way categories are separable from text alone. In particular, Happy–Sad confusion is higher than Happy–Excited confusion in our analysis, and Happy absorbs a substantial portion of Excited instances, yielding an asymmetric confusion pattern. This is consistent with the interpretation that acoustic intensity cues—often crucial for distinguishing Excited from Happy—are absent in text-only settings, so the model defaults to the more frequent positive label.
These observations do not imply that the 6-way taxonomy is intrinsically invalid; rather, they suggest that in text-only ERC, some category boundaries may be weakly identified. This motivates alternative formulations, such as (i) hierarchical classification (e.g., valence first, then finer labels), (ii) dimensional prediction (valence/arousal), or (iii) reduced label sets that merge categories that are difficult to separate without prosody.